For extremely cold temperatures, an automobile manufacturer recommends that a 80% antifreeze and 20% water mix be used in the cooling system of a car. How many quarts of pure (100%) antifreeze and 50% antifreeze and 50% water mix should be combined to make 15 quarts of 80% antifreeze and 20% water mix?
Answer by ankor@dixie-net.com(22740)


For extremely cold temperatures, an automobile manufacturer recommends that a 80% antifreeze and 20% water mix be used in the cooling system of a car.
How many quarts of pure (100%) antifreeze and 50% antifreeze and 50% water mix should be combined to make 15 quarts of 80% antifreeze and 20% water mix?
:
Let x = amt of pure antifreeze
then
(15-x) = amt of 50:50 mix required
:
An amt of antifreeze equation
.50(15-x) + x = .80(15)
7.5 - .5x + 1x = 12
.5x = 12 - 7.5
.5x = 4.5
x = 4.5/.5
x = 9 quarts of pure antifreeze required
: